What do the roots, (-7,0) (-3.0)(1,0)(5,0) of the equations: x^4+4x^3-34x^2-76x+105 tell me besides they are the x intercept?
Found 2 solutions by jim_thompson5910, ewatrrr:
Answer by jim_thompson5910(35256)

They tell you that x^4+4x^3-34x^2-76x+105 factors to (x+7)(x+3)(x-1)(x-5)


Answer by ewatrrr(24785)

 
Hi,
x^4+4x^3-34x^2-76x+105 = (x+7)(x+3)(x-1)(x-5) = 0
roots, (-7,0) (-3.0)(1,0)(5,0)
Knowing the x-intercepts is an important tool in sketching the graph of the function